# Agglomerative Clustering of Simulation Output Distributions Using Regularized Wasserstein Distance

### v1

@@ 3.2 An Agglomerative Clustering Algorithm
Agglomerative clustering is a hierarchical clustering method that begins by treating each instance as an individual cluster and successively merges the closest pairs based on a specified distance metric, allowing clusters to form organically from the data. We choose to employ agglomerative clustering for several reasons. Firstly, unlike $k$-means clustering, in which the number of clusters is predefined, agglomerative clustering excels in situations where the optimal number of clusters is unknown. Secondly, centroid-based methods, such as $k$-means, are sensitive to outliers due to their reliance on a single central point to represent each cluster. Outliers can significantly skew the centroid’s location and distort the clustering process. In contrast, the complete-linkage approach commonly used in agglomerative clustering considers the maximum distance between any two points in distinct clusters, making the clustering more robust to outliers. Complete-linkage clustering also considers the farthest points within the merged clusters, resulting in tighter and more spherical cluster formations compared to single-linkage clustering, which can generate elongated clusters due to chaining effects. Thirdly, agglomerative clustering, particularly when employing complete linkage, offers a valuable output in the form of a dendrogram, which depicts the merging process and the distances between clusters at each stage of the algorithm and can aid in comprehending the relationships between instances and in determining an appropriate number of clusters. Fourthly, unlike the $k$-means algorithm in which one must repeatedly recalculate the centroid of each cluster, agglomerative clustering does not entail calculating centroids. In our setting, the Wasserstein barycenter is a natural choice of centroid, but lacks robustness (Santambrogio & Wang, , 2016); this idiosyncratic behavior of Wasserstein barycenters renders the centroid-based formulation inadequate for representing inter-cluster instances.
@@
To determine the optimal number of clusters, we use the silhouette index proposed by Shahapure & Nicholas, (2020). The silhouette index for a given clustering $\mathcal{C}$ is defined as $S_{\mathcal{C}}=\frac{1}{|\mathcal{C}|}\sum_{C\in\mathcal{C}}\frac{1}{|C|}\sum_{\mu\in C}S_{\mu},$ where $S_{\mu}=(b_{\mu}-a_{\mu})/\max\{b_{\mu},a_{\mu}\}$, $a_{\mu}=(|C_{\mu}|-1)^{-1}\sum_{\mu^{\prime}\in C_{\mu},\,\mu^{\prime}\neq\mu}W_{\lambda}(\mu,\mu^{\prime})$ is the average regularized Wasserstein distance between $\mu$ and every other distribution in the same cluster, and $b_{\mu}=\min_{C^{\prime}\in\mathcal{C},\,C^{\prime}\neq C_{\mu}}\left\{(|C^{\prime}|-1)^{-1}\sum_{\mu^{\prime}\in C^{\prime}}W_{\lambda}(\mu,\mu^{\prime})\right\}$ is the minimum average distance between $\mu$ and distributions in other clusters. The silhouette index considers both intra-cluster (as in $a_{\mu}$) and inter-cluster (as in $b_{\mu}$) distances, and its values fall within the range of $-1$ to $1$, with a higher silhouette index indicating a more favorable clustering. For an individual distribution $\mu$, a silhouette index $S_{\mu}$ close to 1 signifies that $\mu$ is well-positioned within its assigned cluster.
@@
We now present Algorithm 2, an agglomerative algorithm for clustering the multivariate empirical distributions of simulation outputs. In Algorithm 2, $\mathcal{D}$ denotes the distance metric used to calculate the cost matrix between points in the supports. Before applying Algorithm 2, the output data is normalized within each dimension to ensure that no one KPI skews the clustering results.
@@
1:Input: $\{\mu_{1},\ldots,\mu_{N}\}$, $\lambda$, $\mathcal{D}$
@@
To further assess the practicality of Algorithm 2, it is essential to consider its computational cost. Step 1 of Algorithm 2 calculates the pairwise distances between all distributions, the cost of which scales quadratically with the number of distributions, $N$. Calculating the regularized Wasserstein distance between a pair of distributions with the same support size exhibits a quadratic dependence on the size of the support (Altschuler et al., , 2017). After obtaining the pairwise distances, the rest of Algorithm 2 has a cubic dependence on the number of distributions (Karthikeyan et al., , 2020). Additionally, calculating the silhouette index for a given clustering scales quadratically with the number of distributions (Mur et al., , 2016).
@@ 3.3 Wasserstein Barycenter
After clustering the distributions, we turn to the regularized Wasserstein barycenter to summarize the information in each cluster. For a given cluster, the regularized Wasserstein barycenter minimizes the average regularized Wasserstein distance between itself and each of the distributions within the cluster, effectively acting as an “average” of the distributions. To compute the barycenter for a cluster $C$, denoted generically by $\bar{\mu}$, we employ a method that assumes that all distributions have a common support. To conform with this assumption, we manipulate the probability mass vectors of the distributions in each cluster. Specifically, let $\mathcal{Y}_{\bar{\mu}}:=\bigcup_{\mu\in C}\mathcal{Y}_{\mu}$ be the collective support of all distributions in cluster $C$ and let ${M}_{\bar{\mu}}:=|\mathcal{Y}_{\bar{\mu}}|$. For each $\mu\in C$, $\mathbf{p}_{\mu}$ can be modified by extending it to a length of ${M}_{\bar{\mu}}$ by assigning probability masses of 0 to values in $\mathcal{Y}_{\bar{\mu}}\backslash\mathcal{Y}_{\mu}$, resulting in a modified distribution $\tilde{\mu}$ defined on $\mathcal{Y}_{\bar{\mu}}$. The regularized Wasserstein barycenter is a discrete distribution on $\mathcal{Y}_{\bar{\mu}}$ having a probability mass vector
$\mathbf{p}_{\bar{\mu}}:=\underset{\mathbf{p}\in\Delta_{M_{\bar{\mu}}}}{\operatorname{argmin}}\frac{1}{|C|}\sum_{\mu\in C}W_{\lambda}(\tilde{\mu},\bar{\mu}).$
Although the Wasserstein barycenter can be derived by minimizing a weighted sum of regularized Wasserstein distances, in this paper we assume that the scenarios are equally important, and hence weighted equally. The optimal probability mass vector $\mathbf{p}_{\bar{\mu}}$ can be computed using another iterative procedure, given in Algorithm 3.
Algorithm 3 Wasserstein barycenter computation (Benamou et al., , 2015)
1:Input: $C$, $\lambda$, $\mathbf{D}$
2:Construct the matrix $\mathbf{Q}\in\mathbb{R}^{M_{\bar{\mu}}\times M_{\bar{\mu}}}$ having entries $Q_{ll^{\prime}}=e^{-\frac{D_{ll^{\prime}}}{\lambda}}$ for $l,l^{\prime}=1,2,\ldots,M_{\bar{\mu}}$.
3:Initialize vectors $\mathbf{v}_{\mu}^{(0)}=\mathbf{1}_{M_{\bar{\mu}}}$ for all $\mu\in C$.
4:while stopping criteria not met do
@@
6:         Update $\mathbf{u}_{\mu}^{(m)}=\frac{\mathbf{p}_{\bar{\mu}}^{(m)}}{\mathbf{Q}\mathbf{v}_{\mu}^{(m)}}$ and $\mathbf{v}_{\mu}^{(m+1)}=\frac{\mathbf{p}_{\mu}}{\mathbf{Q}^{T}\mathbf{u}_{\mu}^{(m)}}$
@@
8:     Compute the current estimate of the barycenter $\mathbf{p}_{\bar{\mu}}^{(m)}=\prod_{\mu\in C}\left(\mathbf{u}_{\mu}^{(m)}\odot(\mathbf{Q}\mathbf{v}_{\mu}^{(m)})\right)$, where $\Pi$ and $\odot$ denote the element-wise product.
@@
10:Return $\mathbf{p}_{\bar{\mu}}$.
@@ 4.1 Staffing a Fixed Number of Operators
To more deeply understand the distributions within each cluster, we compute the barycenters of each cluster and plot the marginal cumulative distribution functions (cdfs) for each of the five KPIs. In Figure 2, we observe that no cluster consistently outperforms the others, however, Cluster 4 performs well across all five KPIs, whereas the other clusters perform poorly in at least one KPI.
Figure 2: Marginal cdfs for each KPI for each barycenter, for a fixed number of operators.
Having identified a good cluster, we examine the correlation matrix of Cluster 4 in Figure 3. We observe a positive correlation between the overwork of premium service operators and the time spent by both types of customers in the call center, as well as the mean overwork of basic service operators. This suggests that for staffing configurations in Cluster 4, a high overwork time for premium service operators on any given day is associated with both regular and premium customers spending more time in the call center than usual and basic service operators experiencing more overwork time than average. We compare to the correlation matrix of Cluster 5, which performs very well in the first four KPIs but poorly in terms of mean overwork for technical operators. In Figure 3, we observe a strong negative correlation between the mean time in the system of regular and premium customers.
Figure 3: (Left) Correlation matrix for Cluster 4. (Right) Correlation matrix for Cluster 5.
As seen in Figure 4, the staffing configurations that make up Cluster 4 are characterized by having a moderate number of technical operators and a variable number of basic service and premium service operators, where the number of basic service operators is as low as 1 in some configurations. The decision maker might want to identify the staffing configuration among those in Cluster 4 that, say, minimizes the total staffing costs. For instance, if the staffing costs for basic service, premium service, and technical operators were 4, 1, and 1, respectively, then configuration $(7,28,14)$ would be the cheapest.
Figure 4: Simulated staffing configurations, represented as triplets of (# of premium service operators, # of basic service operators, # of technical operators). Configurations in Cluster 4, including the optimal staffing configuration (7, 28, 14), are highlighted.
@@ 4.2 Staffing Subject to a Budget
Staffing-cost considerations could alternatively be incorporated into the design of the experiment. As a follow-up to the staffing-cost analysis above, suppose the decision maker’s objective were to find a staffing configuration with a desirable output distribution among those having total costs between 50 and 55. There are 2,143 feasible staffing configurations and, as before, we uniformly select 100 configurations, simulate each for 40 days, and apply Algorithm 2 on the results. The silhouette index recommends five clusters; the marginal distributions of the corresponding barycenters are shown in Figure 5. Unlike in the previous experiment, no cluster dominates across all five KPIs: each cluster performs very well in at least one KPI, but suffers in other aspects. The decision maker’s priorities play a crucial role in balancing the tradeoffs across KPIs. For instance, if providing good customer service is more important than ensuring favorable conditions for operators, then Cluster 1 is preferable. Conversely, if keeping operators’ overwork low is a priority, then Cluster 3 might be preferred.
Figure 5: Marginal cdfs for each KPI for each barycenter, for staffing subject to a budget.

### v2

@@ 3.2 An Agglomerative Clustering Algorithm
Agglomerative clustering is a hierarchical clustering method that begins by treating each instance as an individual cluster and successively merges the closest pairs based on a specified distance metric, allowing clusters to form organically from the data. We choose to employ agglomerative clustering for several reasons. Firstly, unlike $k$-means clustering, in which the number of clusters is predefined, agglomerative clustering is a better choice in situations where the optimal number of clusters is unknown. Secondly, centroid-based methods, such as $k$-means, are sensitive to outliers due to their reliance on a single central point to represent each cluster. Outliers can significantly skew the centroid’s location and distort the clustering process. In contrast, the complete-linkage approach commonly used in agglomerative clustering considers the maximum distance between any two points in distinct clusters, making the clustering more robust to outliers. Complete-linkage clustering also considers the farthest points within the merged clusters, resulting in tighter and more spherical cluster formations compared to single-linkage clustering, which can generate elongated clusters due to chaining effects. Thirdly, agglomerative clustering, particularly when employing complete linkage, offers a valuable output in the form of a dendrogram, which depicts the merging process and the distances between clusters at each stage of the algorithm and can aid in comprehending the relationships between instances and in determining an appropriate number of clusters. Fourthly, unlike the $k$-means algorithm in which one must repeatedly recalculate the centroid of each cluster, agglomerative clustering does not entail calculating centroids.
To determine the optimal number of clusters, we use the silhouette index proposed by Shahapure and Nicholas (2020). The silhouette index for a given clustering $\mathcal{C}$ is defined as
$S_{\mathcal{C}}=\frac{1}{|\mathcal{C}|}\sum_{C\in\mathcal{C}}\frac{1}{|C|}\sum_{\mu\in C}S_{\mu},$
where $S_{\mu}=(b_{\mu}-a_{\mu})/\max\{b_{\mu},a_{\mu}\}$,
$a_{\mu}=(|C_{\mu}|-1)^{-1}\sum_{\mu^{\prime}\in C_{\mu},\,\mu^{\prime}\neq\mu}W_{\lambda}(\mu,\mu^{\prime})$
is the average regularized Wasserstein distance between $\mu$ and every other distribution in the same cluster, and
$b_{\mu}=\min_{C^{\prime}\in\mathcal{C},\,C^{\prime}\neq C_{\mu}}\left\{(|C^{\prime}|-1)^{-1}\sum_{\mu^{\prime}\in C^{\prime}}W_{\lambda}(\mu,\mu^{\prime})\right\}$
@@
is the minimum average distance between $\mu$ and distributions in other clusters. The silhouette index considers both intra-cluster (as in $a_{\mu}$) and inter-cluster (as in $b_{\mu}$) distances, and its values fall within the range of $-1$ to $1$, with a higher silhouette index indicating a more favorable clustering. For an individual distribution $\mu$, a silhouette index $S_{\mu}$ close to 1 signifies that $\mu$ is well positioned within its assigned cluster.
@@
We now present Algorithm 2, an agglomerative algorithm for clustering the multivariate empirical distributions of simulation outputs. In Algorithm 2, $\mathbf{D}$ denotes the cost matrix between points in the supports. Before applying Algorithm 2, the output data is normalized within each dimension to ensure that no one KPI skews the clustering results.
@@
1:Input: $\{\mu_{1},\ldots,\mu_{N}\}$, $\lambda$, $\mathbf{D}$
@@
To further assess the practicality of Algorithm 2, it is essential to consider its computational cost. Step 1 of Algorithm 2 calculates the pairwise distances between all distributions, the cost of which scales quadratically with the number of distributions, $N$. Calculating the regularized Wasserstein distance between a pair of distributions with the same support size exhibits a quadratic dependence on the size of the support (Altschuler et al., 2017). After obtaining the pairwise distances, the rest of Algorithm 2 has a cubic dependence on $N$ (Karthikeyan et al., 2020). Additionally, calculating the silhouette index for a given clustering scales quadratically with $N$ (Mur et al., 2016).
Remark In experiments with stochastic simulation models, the user has complete control over how pseudo-random numbers are used to drive the simulation model. One such scheme is common random numbers (CRN), which refers to the technique of using the same streams of pseudo-random numbers when simulating different scenarios in an attempt to induce positive correlation in the outputs on any given replication (Kelton, 2006). When successful, CRN enables more accurate comparisons between scenarios and is a powerful tool for simulation optimization. We believe that CRN can have benefits for clustering as well. In particular, we conjecture that CRN will tend to reduce the variance in the Wasserstein distances between output distributions and consequently reduce the variance of the clustering outcomes. Furthermore, we anticipate that CRN should yield clusterings that more closely resemble the true underlying clustering structure, i.e., how the scenarios’ true outputs distributions would be clustered were they known. As we do not believe these benefits of CRN are universal, we eschew attempting to prove theoretical statements about them and instead provide some empirical support for these assertions in Section 4.
@@ 3.3 Wasserstein Barycenter
After clustering the distributions, we turn to the regularized Wasserstein barycenter to summarize the information in each cluster. For a given cluster, the regularized Wasserstein barycenter minimizes the average regularized Wasserstein distance between itself and each of the distributions within the cluster, effectively acting as an “average” of the distributions. There are several approaches to compute the barycenter for a cluster $C$, denoted generically by $\bar{\mu}$. One approach is to fix the support and then optimize over the probabilities assigned to each support point. Alternatively, we could fix the number of support points and their corresponding probabilities, as discussed in Altschuler et al. (2017), and then optimize over the locations of the support points. We choose to employ an iterative algorithm introduced in Cuturi and Doucet (2014) to simultaneously optimize both the support points and their corresponding probabilities, with the only fixed parameter being the support size. This procedure continues until the stopping criterion, defined similarly to Algorithm 1, is met. Specifically we let the support of the barycenter be $\mathcal{Y}_{\bar{\mu}}$. The regularized Wasserstein barycenter of a cluster $C$ is a discrete distribution on $\mathcal{Y}_{\bar{\mu}}$ having a probability mass vector
$\mathbf{p}_{\bar{\mu}}:=\underset{\mathbf{p}\in\Delta_{M_{\bar{\mu}}}}{\operatorname{argmin}}\frac{1}{|C|}\sum_{\mu\in C}W_{\lambda}(\mu,\bar{\mu}).$
In this paper, we assume that the scenarios are equally important, and hence equally weighted, but the Wasserstein barycenter can be derived by minimizing a weighted sum of regularized Wasserstein distances The optimal probability mass vector $\mathbf{p}_{\bar{\mu}}$ for a fixed support can be computed using another iterative procedure, given in Algorithm 3. Then by incorporating Algorithm 3 into Algorithm 4, we can allow the support to be free. Although we have regarded the supports as sets, in Algorithm alg: reg barycenter we will denote the supports $\mathcal{Y}_{\bar{\mu}}$ and $\mathcal{Y}_{\mu}$ as matrices $\boldsymbol{Y_{\bar{\mu}}}$ and $\boldsymbol{Y_{\mu}}$, with the understanding that the ordering of elements within these matrices is arbitrary. The parameters $\beta$ and $\theta$ appearing in Algorithms 3 and 4 vary between 0 and 1 and control how the step sizes are updated.
Algorithm 3 Finding the optimal $p_{\bar{\mu}}$ for a fixed support (Cuturi and Doucet, 2014)
1:Input: C, $\lambda$, $\mathbf{D}$, $\beta$
2:Set $\hat{p}_{\bar{\mu}}=\tilde{p}_{\bar{\mu}}=\mathbf{1}/{|\mathcal{Y}_{\bar{\mu}}|}$.
3:while not converged do
4:     $\beta=(t+1)/2$, $p_{\bar{\mu}}=(1-\beta^{-1})\hat{p}_{\bar{\mu}}+\beta^{-1}\tilde{p}_{\bar{\mu}}$.
5:     Calculate $\alpha_{\bar{\mu}}=\frac{1}{|C|}\sum_{\mu\in C}\alpha_{\mu}$ where $\alpha_{\mu}=-\lambda\log(u_{\mu})$ is derived when finding the regularized Wasserstein distance between $\mu$ and $\bar{\mu}$ using Algorithm 1.
6:     $\tilde{p}_{\bar{\mu}}\leftarrow\tilde{p}_{\bar{\mu}}\odot e^{-t_{0}\beta\alpha_{\bar{\mu}}}$, $\tilde{p}_{\bar{\mu}}\leftarrow\frac{\tilde{p}_{\bar{\mu}}}{\tilde{p}_{\bar{\mu}}^{T}\mathbf{1}_{|C|}}$.
7:     $\hat{p}_{\bar{\mu}}\leftarrow(1-\beta^{-1})\hat{p}_{\bar{\mu}}+\beta^{-1}\tilde{p}_{\bar{\mu}}$, $t\leftarrow t+1$.
8:end while
9:return $p_{\bar{\mu}}$.
Algorithm 4 Regularized Wasserstein barycenter (Cuturi and Doucet, 2014)
1:Input: C, $\lambda$, $\mathbf{D}$, $\beta$, $\theta$.
2:Initialize $\boldsymbol{Y}_{\bar{\mu}}$ and $p_{\bar{\mu}}$.
3:while $\mathcal{Y}_{\bar{\mu}}$ and $p_{\bar{\mu}}$ have not converged do
4:     Derive $p_{\bar{\mu}}$ using Algorithm 3.
@@
6:         Obtain $\gamma_{\lambda,\mu}$ using $W_{\lambda}(\bar{\mu},\mu)$.
@@
8:     $\boldsymbol{Y}_{\bar{\mu}}=(1-\theta)\boldsymbol{Y}_{\bar{\mu}}+\theta\left(\frac{1}{|C|}\sum_{\mu\in C}\boldsymbol{Y}_{\mu}\gamma_{\lambda,\mu}\right)\text{diag}(p_{\bar{\mu}})^{-1}$.
@@
10:return $\boldsymbol{Y}_{\bar{\mu}}$, $p_{\bar{\mu}}$.
The Wasserstein barycenter lacks robustness (Santambrogio and Wang, 2016), and this idiosyncratic behavior makes centroid-based clustering approaches like $k$-means inadequate for representing inter-cluster variability. Additionally, agglomerative clustering is more efficient in terms of computational cost. After introducing the barycenter concept, we compare the efficiency of our hierarchical algorithm with that of a $k$-means algorithm using barycenters defined in Algorithm 4 for the cluster centroids. In Figure 2, we fix the number of scenarios under consideration and vary the number of replications from 10 to 300, common across scenarios. In a second experiment, we fix the support size—the number of replications—and vary the number of scenarios from 10 to 300. For the $k$-means algorithm, we do not know the optimal number of clusters and therefore we choose to calculate the clusterings for 2 to 10 clusters. In Figure 2, we observe that the computational cost increases with the number of scenarios for both $k$-means and agglomerative clustering, however, the rate of increase is significantly smaller for agglomerative clustering. The relatively high computational cost of the $k$-means algorithm is due to the repeated calculation of barycenters. The number of such calculations depends on the number of systems and clusters, whereas the time for each calculation depends on the support size of the distributions and the barycenter. In this experiment, the support size of the barycenter is taken to be the same as that of the input distributions.
Figure 2: Comparison of computational times for $k$-means and agglomerative clustering for varying numbers of distributions with a fixed support size of 20 (Left) and for varying support sizes for 20 distributions (Right). The variability was negligible compared to the difference in computational costs between the two algorithms.
@@ 4.1 Staffing a Fixed Number of Operators
To more deeply understand the distributions within each cluster, we compute the barycenter of each cluster and plot the kernel density estimates of probability density functions (pdfs) for each of the five KPIs. In Figure 4, we observe that no cluster consistently outperforms the others; however, Cluster 4 performs well across all five KPIs, whereas the other clusters perform poorly in at least one KPI.
Figure 4: Marginal pdfs for each KPI for each barycenter, for a fixed number of operators.
Another cluster of interest is Cluster 1, as it performs well in two KPIs and moderately (relative to other clusters) in the remaining three KPIs. To understand this behavior better, we can examine the correlation matrices of the distributions in this cluster, depicted in Figure 5. The matrices reveal strong positive correlations between the mean regular customer waiting time, $Y_{1}$, the overwork of basic operators, $Y_{3}$, and the overwork of premium operators, $Y_{4}$. This suggests that on days for which the regular customers’ waiting time is above average, the workload increased for both types of operators, and on any given day observing one of these correlated metrics can potentially reveal the trend of the others.
Figure 5: Correlation matrices for staffing configurations in Cluster 1.
Returning to our analysis of Cluster 4, it can be seen in Figure 6 that the staffing configurations that comprise Cluster 4 are characterized by having a moderate number of premium and technical operators and a variable number of basic service operators ranging from 1 to 30. The decision maker might want to identify the staffing configuration in Cluster 4 that, say, minimizes the total staffing costs. For instance, if the staffing costs for basic service, premium service, and technical operators were 4, 1, and 1, respectively, then the configuration with 7 specialist, 28 regular, and 14 technical operators, would be the cheapest. Rather than choose staffing configuration (7, 28, 14) on the basis of cost alone, we recommend further inspecting the individual distributions within Cluster 4, as barycenters predominantly capture central tendencies and may conceal significant intra-cluster variability or outliers. The marginal cdfs of the configurations in Cluster 4, including that of its barycenter, are illustrated in Figure 7. Upon further evaluation, we observe that staffing configuration $(7,28,14)$ performs adequately across $Y_{1}$, $Y_{4}$, and $Y_{5}$, however, its performance in $Y_{2}$ and $Y_{3}$ is unsatisfactory. If we were to sort the staffing configurations in Cluster 4 by their costs and continue down the list, we would find that the second cheapest configuration in Cluster 4, (9, 23, 17), demonstrates solid performance across all metrics. However, staffing configuration (9, 23, 17) slightly underperforms the barycenter with respect to $Y_{1}$. If we were to insist on identifying a staffing configuration that stochastically dominates the barycenter in every performance, which many not necessarily exist, we would find that the sixth cheapest configuration, (12, 18, 19), is the cheapest configuration that does so.
Figure 6: Simulated staffing configurations, represented as triplets of (# of premium service operators, # of basic service operators, # of technical operators). Configurations in Cluster 4, including the cheapest (7, 28, 14), second cheapest (9, 23, 17), and sixth cheapest (12, 18, 19) configurations in Cluster 4, are highlighted.
Figure 7: Comparison of marginal cdfs for different staffing configurations in Cluster 4.
@@ 4.2 Staffing Subject to a Budget
Staffing-cost considerations could alternatively be incorporated into the design of the experiment. As a follow-up to the staffing-cost analysis above, suppose the decision maker’s objective were to find a staffing configuration with a desirable output distribution among those having total costs between 50 and 55, assuming the same per-operator costs as before. There are 2,143 feasible staffing configurations, and we again uniformly select 100 configurations, simulate each for 40 days, and apply Algorithm 2 on the results. The silhouette index recommends five clusters, and smoothed versions of the marginal pdfs of the corresponding barycenters are shown in Figure 8. Unlike in the previous experiment, no cluster dominates across all five KPIs: each cluster performs well in at least one KPI, but suffers in other aspects. The decision maker’s priorities thus play a crucial role in balancing the tradeoffs across KPIs. For instance, if providing good customer service is more important than ensuring favorable conditions for operators, then Cluster 1 is preferable. Conversely, if keeping basic or technical operators’ overwork low is a priority, then Cluster 3 might be preferred. Further investigation by the decision maker can identify the most cost-effective option or the configuration that outperforms other configurations in certain KPIs.
Figure 8: Marginal pdfs for each KPI for each barycenter for staffing subject to a budget.